At a health food store, dried apple slices that cost $1.80 per pound are mixed with dried banana slices that cost $2.10 per pound. If the mixture weighs 5 pounds and costs $1.92 per pound, how many pounds of apple slices were used?
A) 1.5 pounds of apples B) 2 pounds of apples C) 3 pounds of apples
D) 3.5 pounds of apples E) none of these

Answer by checkley77(12569)

2.10X+1.80(5-X)=1.92*5
2.10X+9-1.80X=9.6
.3X=9.6-9
.3X=.6
X=.6/.3
X=2 POUNDS OF BANANA SLICES IS USED.
5-2=3 POUNDS OF APPLE SLICES IS USED.
ANSWER C).
PROOF:
2.10*2+1.8*3=1.92*5
4.2+5.4=9.6
9.6=9.6